# Absolute continuity of non-Gaussian and Gaussian multiplicative chaos measures

### v1

@@ 3.2.1. Reduction, and overview of the proof
Recall the notation introduced in (3.23). A union bound over the set of generations at which the field first becomes more than $\alpha$ thick simply yields
$M_{n,\widetilde{a}}^{(2)}(w)\leq\lim\limits_{r\to\infty}\sum_{l=\log(n)^{2}}^{r}|I_{n}(w)|^{-1}\int_{I_{n}(w)}\mathbbm{1}_{A_{l;n}(t)}e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)^{-1}\;\mathrm{d}t\,.$
Using generational independence and (2.9), we have $\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)=e^{\log(2)\frac{\gamma^{2}}{2}l+O_{r}(1)}\widetilde{Z}_{(n+l,n+r],\gamma,\widetilde{a}}(t)$. Moreover, by definition, we have
$\mathbbm{1}_{A_{l,n}(t)}\leq e^{(\alpha-\gamma)\widetilde{S}_{l,\widetilde{a}}(t)-(\alpha-\gamma)\alpha\log(2)l}\,.$
Substituting into the second-to-last display then applying Fatou, Fubini, and Lemma 2.2 yield
$\mathbb{E}\left[M_{n,\widetilde{a}}^{(2)}(w)\right]\leq C\sum_{l=\log(n)^{2}}^{\infty}2^{\frac{\alpha^{2}}{2}l-(\alpha-\gamma)\alpha l-\frac{\gamma^{2}}{2}l}\leq C2^{-\frac{(\alpha-\gamma)^{2}}{2}\log(n)^{2}}\,.\qed$
@@
First, we control the $\mathbb{E}[M_{n,\widetilde{a}}^{(1)}(w_{1})M_{n,\widetilde{a}}^{(1)}(w_{2})]$ term. For this we set $D:=I_{n}(w_{1})\times I_{n}(w_{2})$. Using Fatou and Fubini, and bounding indicators by $1$, we have
$\mathbb{E}[M_{n,\widetilde{a}}^{(1)}(w_{1})M_{n,\widetilde{a}}^{(1)}(w_{2})]$
$\leq\liminf\limits_{r\to\infty}|D|^{-1}\int_{D}\mathbb{E}\left[e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)+\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(s)}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)^{-1}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(s)^{-1}\right]\mathrm{d}t\mathrm{d}s\,.$ (3.27)
By Lemma 2.2 and the definition of $\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}$ we have
@@
Similar to the calculation in [Junnila, Section 3.1, pp. 6190,6191] we get for $(t,s)\in D$ (recalling that $v_{n}(t)=w_{1}$, $v_{n}(s)=w_{2}$),
@@
where we used the assumption that $|t_{n}(w_{1})-t_{n}(w_{2})|\geq 2^{-n}$, which is much larger than $\max(|t-t_{n}(w_{1})|,|s-t_{n}(w_{2})|)$ in light of (2.1). Plugging (3.28) and (3.29) into (3.2.3) yields
@@
$\mathbb{E}\Big{[}M_{n,\widetilde{a}}^{(1)}(w_{1})\Big{]}\mathbb{E}\Big{[}M_{n,\widetilde{a}}^{(1)}(w_{2})\Big{]}\geq\Big{(}1-C2^{-\frac{(\alpha-\gamma)^{2}}{2}\log(n)^{2}}\Big{)}^{2}\geq 1-C2^{-\frac{(\alpha-\gamma)^{2}}{2}\log(n)^{2}}\,.$
Thus
@@ 3.3. Proof of (C): uniform integrability
Similarly to the last section we only show that $(\widetilde{R}_{n,\gamma}(t))_{n\in\mathbb{N}}$ is almost-surely uniformly integrable with respect to $\widetilde{\mu}_{\gamma,\widetilde{a}}$, since the other statement can be shown analogously. By the coupling construction, we can control $\widetilde{R}_{n,\gamma}(t)$ for all $t$ which are “consistently thick”, with an error depending on how late they start being consistently thick. Thus, conceptually we aim to show that points which only start being consistently $(\gamma-\delta)\log(2)$-thick at a late time don’t contribute to $\widetilde{\mu}_{\gamma,\widetilde{a}}$. We split the calculation into the $t$ which are at least $(\gamma-\delta)\log(2)$-thick at time $n$ and those who aren’t. Recall $\mathcal{T}_{n,\gamma}$ from (2.19). Say that $t\in\mathcal{T}_{n,\gamma}$ if and only if $v_{n}(t)\in\mathcal{T}_{n,\gamma}$, and set
@@
Since $\widetilde{R}_{n,\gamma}(t)=\widetilde{R}_{n,\gamma,1}(t)+\widetilde{R}_{n,\gamma,2}(t)$, we are done upon showing $(\widetilde{R}_{n,\gamma,1}(t))_{n\in\mathbb{N}_{0}}$, $(\widetilde{R}_{n,\gamma,2}(t))_{n\in\mathbb{N}}$ are both uniformly integrable with respect to $\widetilde{\mu}_{\gamma,\widetilde{a}}$.
@@
$J_{\eta,n}$ $=\sum_{v\in\mathcal{N}_{n}}\widetilde{\mu}_{\gamma,\widetilde{a}}[I_{n}(v)]\widetilde{R}_{n,\gamma,1}(t_{n}(v))^{1+\eta}\leq C\sum_{v\in\mathcal{N}_{n}}\mathbbm{1}_{\{v\in\mathcal{T}_{n,\gamma}\}}\widetilde{\mu}_{\gamma,\widetilde{a}}[I_{n}(v)]e^{\gamma(1+\eta)(\widetilde{S}_{n,\widetilde{g}}(v)-\widetilde{S}_{n,\widetilde{a}}(v))},$
where in the second step, we used Eq. 2.9 to bound both $\widetilde{Z}_{n,\gamma,\widetilde{a}}(t)$ and $\widetilde{Z}_{n,\gamma,\widetilde{g}}(t)$ by $Ce^{\log(2)\frac{\gamma^{2}}{2}n}$.
Now we decompose the sum according to the time $L\in\{1,\dots,n\}$ at which $v$ starts to be consistently thick, i.e., we set
$\mathcal{T}_{L,n,\gamma}:=\bigg{\{}v\in\mathcal{N}_{n}:\frac{\max\{\widetilde{S}_{L-1,\widetilde{a}}(v),\widetilde{S}_{L-1,\widetilde{g}}(v)\}}{\log(2)(L-1)}\leq\gamma-\delta,\forall_{k\in\{L,\dots,n\}}\,\frac{\max\{\widetilde{S}_{L,\widetilde{a}}(v),\widetilde{S}_{L,\widetilde{g}}(v)\}}{\log(2)k}>\gamma-\delta\bigg{\}}\,$
[…]
@@
which is $\mathbb{P}$-almost surely finite since $L^{\ast}$ and $\mathcal{N}_{L^{\ast}}$ are. Thus it suffices to prove that, $\mathbb{P}$-almost surely,
$\widetilde{J}_{\eta,n}:=\sum_{L=L^{\ast}+1}^{n}\sum_{v\in\mathcal{T}_{L,n,\gamma}}\widetilde{\mu}_{\gamma,\widetilde{a}}[I_{n}(v)]e^{\gamma(1+\eta)(\widetilde{S}_{L,\widetilde{g}}(v)-\widetilde{S}_{L,\widetilde{a}}(v))}<\infty\,.$
@@
and use that $\widetilde{Z}_{L,\gamma,\widetilde{a}}(w)\leq C2^{\frac{\gamma^{2}}{2}\log(L)}$ by (2.9) to obtain
$\widetilde{J}_{\eta,n}\leq C\sum_{L=L^{\ast}+1}^{n}\sum_{w\in\mathcal{N}_{L}}\mathbbm{1}_{\{w\in\mathcal{T}_{L,\gamma}\}}|I_{L}(w)|\,M_{L,\widetilde{a}}(w)e^{-\gamma\eta\widetilde{S}_{L,\widetilde{a}}(w)+\gamma(1+\eta)\widetilde{g}_{L}(w)}2^{\gamma(1+\eta)(\gamma-\delta)L-\frac{\gamma^{2}}{2}L}.$ (3.32)
[…]
@@
is $\mathbb{P}$-almost surely finite. Substituting (3.33), (3.34) into (3.32) yields that there is a $\mathbb{P}$-a.s. random variable $\mathcal{C}$ such that
$\widetilde{J}_{\eta,n}\leq\mathcal{C}\sum_{L=L^{\ast}+1}^{n}2^{\gamma(1+\eta)(\gamma-\delta)L+\varepsilon L+\mathcal{C}_{1}L^{2/3}-\frac{\gamma^{2}}{2}L}2^{-L}f(L)^{-1}\sum_{w\in\mathcal{N}_{L}}\mathbbm{1}_{\{w\in\mathcal{T}_{L,\gamma}\}}M_{L,\widetilde{a}}(w)=:\mathcal{C}\sum_{L=L^{\ast}+1}^{n}\sigma_{L},$
@@
For $\eta=\varepsilon=0$ this is bounded by $O_{L}(L^{-2}2^{-\delta^{2}L}2^{\mathcal{C}_{1}L^{2/3}})\leq O_{L}(2^{-\varepsilon^{\prime}L})$ for some $\varepsilon^{\prime}>0$. Since the expression is continuous in both $\eta$ and $\varepsilon$, we can choose them so small that $\mathbb{E}[\sigma_{L}]\leq O_{L}(2^{-\varepsilon^{\prime}/2L})$. By Markov’s inequality and Borel-Cantelli, this implies that for all but finitely many $L$ we have $\sigma_{L}\leq 2^{-\varepsilon^{\prime}/4L}$. This immediately implies that $\mathbb{P}$-almost surely we have
$\widetilde{J}_{\eta,n}<\infty\,,$
finishing the proof of absolute continuity of $\widetilde{R}_{n,\gamma,1}(t)$ with respect to $\widetilde{\mu}_{\gamma,\widetilde{a}}$.
@@
$J_{\eta,n}^{\prime}:=\sum_{v\in\mathcal{N}_{n}}\mathbbm{1}_{\{t\in\mathcal{T}_{n,\gamma}^{c}\}}\widetilde{\mu}_{\gamma,\widetilde{a}}[I_{n}(v)]e^{\gamma(1+\eta)(\widetilde{S}_{n,\widetilde{g}}(v)-\widetilde{S}_{n,\widetilde{a}}(v))}\,.$
Similar to Eq. 3.32, we have
@@
[…]
@@
[…]

### v2

@@ 3.2.1. Reduction, and overview of the proof
Recall the notation introduced in Eq.˜3.23. A union bound over the set of generations at which the field first becomes more than $\alpha$ thick simply yields
$M_{n,\widetilde{a}}^{(2)}(w)\leq\lim\limits_{r\to\infty}\sum_{\ell=\log(n)^{2}}^{r}|I_{n}(w)|^{-1}\int_{I_{n}(w)}\mathbbm{1}_{A_{\ell;n}^{\alpha}(t)}e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)^{-1}\;\mathrm{d}t\,.$
By generational independence and Eq.˜2.9, we get
$\widetilde{Z}_{(n,n+r],\thinspace\gamma,\widetilde{a}}(t)\leq Ce^{\log(2)\frac{\gamma^{2}}{2}l}\widetilde{Z}_{(n+\ell,n+r],\gamma,\widetilde{a}}(t)\,.$
Moreover, by definition, we have
$\mathbbm{1}_{A_{\ell;n}^{\alpha}(t)}\leq e^{(\alpha-\gamma)\widetilde{S}_{\ell,\widetilde{a}}(t)-(\alpha-\gamma)\alpha\log(2)\ell}\,.$
Substituting into the second-to-last display then applying Fatou’s lemma, Fubini’s theorem, and Lemma˜2.2 yield
$\mathbb{E}\left[M_{n,\widetilde{a}}^{(2)}(w)\right]\leq C\sum_{\ell=\log(n)^{2}}^{\infty}2^{\frac{\alpha^{2}}{2}\ell-(\alpha-\gamma)\alpha\ell-\frac{\gamma^{2}}{2}\ell}\leq C2^{-\frac{(\alpha-\gamma)^{2}}{2}\log(n)^{2}}\,.\qed$
@@
First, we control the $\mathbb{E}[M_{n,\widetilde{a}}^{(1)}(w_{1})M_{n,\widetilde{a}}^{(1)}(w_{2})]$ term. For this we set $D:=I_{n}(w_{1})\times I_{n}(w_{2})$. Using Fatou’s lemma and Fubini’s theorem, and bounding indicators by $1$, we have
$\mathbb{E}[M_{n,\widetilde{a}}^{(1)}(w_{1})M_{n,\widetilde{a}}^{(1)}(w_{2})]$ (3.27)
$\leq\liminf\limits_{r\to\infty}|D|^{-1}\int_{D}\mathbb{E}\left[e^{\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(t)+\gamma\widetilde{S}_{(n,n+r],\widetilde{a}}(s)}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(t)^{-1}\widetilde{Z}_{(n,n+r],\gamma,\widetilde{a}}(s)^{-1}\right]\mathrm{d}t\mathrm{d}s\,.$
We now claim that Lemma 2.2 implies
@@
To see how Lemma 2.2 applies here, recall that $\widetilde{S}_{[n,n+r],\widetilde{a}}(t)=\widetilde{S}_{n+r,\widetilde{a}}(t)-\widetilde{S}_{n,\widetilde{a}}(t)$. Re-defining $\varphi(\gamma_{1},\gamma_{2})$ in the fourth display of the proof of Lemma 2.2 so that the lowest index of $k$ is $2^{n}$ instead of $1$ and the highest index of $k$ is $2^{n+r}-1$ instead of $2^{n}-1$, we have $\varphi(\gamma,\gamma)$ equals the left-hand side of (3.28). Following the remainder of the proof reveals that the resulting $\widetilde{\zeta}(\gamma,\gamma,t,s)$ term is bounded by $O_{n}(2^{-n/2})$ uniformly in all parameters.
Following the calculation in [25, Section 3.1, Proof of Theorem 1] until the bound $|R_{n}(t)|\leq\frac{1}{(n+1)\sin(2\pi t)}$ obtained in the last line of that proof, we similarly have (where their $n$, $t$, and $X_{k}(x)$ correspond to $2^{n+h}-1$, tst-s, and $X_{k}(t)$ respectively for us) for $(t,s)\in D$ (and recalling that $v_{n}(t)=w_{1}$, $v_{n}(s)=w_{2}$),
@@
where we used the assumption that $|t_{n}(w_{1})-t_{n}(w_{2})|\geq 2^{-n}$, which is much larger than $\max(|t-t_{n}(w_{1})|,|s-t_{n}(w_{2})|)$ in light of Eq.˜2.1. Plugging Eq.˜3.28 and Eq.˜3.29 into Eq.˜3.27 yields
@@
$\mathbb{E}\Big[M_{n,\widetilde{a}}^{(1)}(w_{1})\Big]\mathbb{E}\Big[M_{n,\widetilde{a}}^{(1)}(w_{2})\Big]\geq\Big(1-C2^{-\frac{(\alpha-\gamma)^{2}}{2}\log(n)^{2}}\Big)^{2}\geq 1-2C\cdot 2^{-\frac{(\alpha-\gamma)^{2}}{2}\log(n)^{2}}\,.$
Thus, absorbing the factor $2$ into the generic constant $C$,
@@ 3.3. Proof of (C): uniform integrability
Similarly to the last section we only show that $(\widetilde{R}_{n,\gamma}(t))_{n\in\mathbb{N}}$ is almost surely uniformly integrable with respect to $\widetilde{\mu}_{\gamma,\widetilde{a}}$, since the other statement can be shown analogously. By the coupling construction, we can control $\widetilde{R}_{n,\gamma}(t)$ for all $t$ which are “consistently thick”, with an error depending on how late they start being consistently thick. Thus, conceptually we aim to show that points which only start being consistently $(\gamma-\delta)\log(2)$-thick at a late level do not contribute to $\widetilde{\mu}_{\gamma,\widetilde{a}}$. We split the calculation into the $t$ which are at least $(\gamma-\delta)\log(2)$-thick at level $n$ and those who are not. Recall $\mathcal{T}_{n,\gamma}$ from (2.19). Say that $t\in\mathcal{T}_{n,\gamma}$ if and only if $v_{n}(t)\in\mathcal{T}_{n,\gamma}$, and set
@@
Since $\widetilde{R}_{n,\gamma}(t)=\widetilde{R}_{n,\gamma,1}(t)+\widetilde{R}_{n,\gamma,2}(t)$, we are done upon showing that $(\widetilde{R}_{n,\gamma,1}(t))_{n\in\mathbb{N}_{0}}$, $(\widetilde{R}_{n,\gamma,2}(t))_{n\in\mathbb{N}}$ are both uniformly integrable with respect to $\widetilde{\mu}_{\gamma,\widetilde{a}}$.
@@
$J_{\eta,n}$ $=\sum_{v\in\mathcal{N}_{n}}\widetilde{\mu}_{\gamma,\widetilde{a}}[I_{n}(v)]\widetilde{R}_{n,\gamma,1}(t_{n}(v))^{1+\eta}$
$\leq C\sum_{v\in\mathcal{N}_{n}}\mathbbm{1}_{\{v\in\mathcal{T}_{n,\gamma}\}}\widetilde{\mu}_{\gamma,\widetilde{a}}[I_{n}(v)]e^{\gamma(1+\eta)(\widetilde{S}_{n,\widetilde{g}}(v)-\widetilde{S}_{n,\widetilde{a}}(v))},$
where in the second step, we used Eq.˜2.9 to bound both $\widetilde{Z}_{n,\gamma,\widetilde{a}}(t)$ and $\widetilde{Z}_{n,\gamma,\widetilde{g}}(t)$ by $Ce^{\log(2)\frac{\gamma^{2}}{2}n}$.
Now we decompose the sum according to the level $L\in\{1,\dots,n\}$ at which $v$ starts to be consistently thick, i.e., we set
$\mathcal{T}_{L,n,\gamma}:=\bigg\{v\in\mathcal{N}_{n}:\frac{\max\{\widetilde{S}_{L-1,\widetilde{a}}(v),\widetilde{S}_{L-1,\widetilde{g}}(v)\}}{\log(2)(L-1)}\leq\gamma-\delta,\\ \forall_{k\in\{L,\dots,n\}}\,\frac{\max\{\widetilde{S}_{L,\widetilde{a}}(v),\widetilde{S}_{L,\widetilde{g}}(v)\}}{\log(2)k}>\gamma-\delta\bigg\}\,$
and note that $\bigcup_{L=1}^{n}\mathcal{T}_{L,n,\gamma}=\mathcal{T}_{n,\gamma}$. By Lemma 2.6 and Borel-Cantelli, there exist $L^{\ast}<\infty$ $\mathbb{P}$-almost surely such that $\widetilde{g}_{r}(v)-\widetilde{a}_{r}(v)\leq r^{-2}$ for all $v\in\mathcal{T}_{L,n,\gamma}$ and $r\geq\max\{L^{*},L+1\}$. Towards Eq.˜3.31, we from now on fix $\omega\in\Omega^{\ast}$. We then have
@@
which is $\mathbb{P}$-almost surely finite since $L^{\ast}$ and $\mathcal{N}_{L^{\ast}}$ are. Thus, it suffices to prove
$\widetilde{J}_{\eta,n}:=\sum_{L=L^{\ast}+1}^{n}\sum_{v\in\mathcal{T}_{L,n,\gamma}}\widetilde{\mu}_{\gamma,\widetilde{a}}[I_{n}(v)]e^{\gamma(1+\eta)(\widetilde{S}_{L,\widetilde{g}}(v)-\widetilde{S}_{L,\widetilde{a}}(v))}<\infty\quad\text{$\mathbb{P}$-almost surely.}$ (3.32)
@@
and use that $\widetilde{Z}_{L,\gamma,\widetilde{a}}(w)\leq C2^{\frac{\gamma^{2}}{2}\log(L)}$ by Eq.˜2.9 to obtain
$\widetilde{J}_{\eta,n}\leq\\ C\sum_{L=L^{\ast}+1}^{n}\sum_{w\in\mathcal{N}_{L}}\mathbbm{1}_{\{w\in\mathcal{T}_{L,\gamma}\}}|I_{L}(w)|\,M_{L,\widetilde{a}}(w)e^{-\gamma\eta\widetilde{S}_{L,\widetilde{a}}(w)+\gamma(1+\eta)\widetilde{g}_{L}(w)}2^{\gamma(1+\eta)(\gamma-\delta)L-\frac{\gamma^{2}}{2}L}.$ (3.33)
Using a union bound and Borel-Cantelli, there exist $c^{\prime}>0$ and a random variable $L^{\ast\ast}\in(0,\infty)$ $\mathbb{P}$-almost surely such that for all $L\geq L^{\ast\ast}$ and $w\in\mathcal{N}_{L}$, we have $\widetilde{S}_{L,\widetilde{a}}(w)\geq-c^{\prime}L$. In particular, there exists $\mathcal{C}(\eta)$ taking values in $(0,\infty)$ $\mathbb{P}$-almost surely such that $\gamma\eta(\widetilde{S}_{L,\widetilde{a}}(w)+c^{\prime}L)\geq\mathcal{C}(\eta)$ for all $L\in\mathbb{N}$. Then for any $\varepsilon>0$, we can choose $\eta$ small such that for all $n\in\mathbb{N}$, $L\leq n$, $w\in\mathcal{T}_{L,\gamma}$,
@@
is $\mathbb{P}$-almost surely finite. Substituting Eq.˜3.34, Eq.˜3.35 into Eq.˜3.33 yields that there is a $\mathbb{P}$-a.s. random variable $\mathcal{C}$ such that
$\widetilde{J}_{\eta,n}$ $\leq\mathcal{C}\sum_{L=L^{\ast}+1}^{n}2^{\gamma(1+\eta)(\gamma-\delta)L+\varepsilon L+\mathcal{C}_{1}L^{2/3}-\frac{\gamma^{2}}{2}L}2^{-L}f(L)^{-1}\sum_{w\in\mathcal{N}_{L}}\mathbbm{1}_{\{w\in\mathcal{T}_{L,\gamma}\}}M_{L,\widetilde{a}}(w)$
$=:\mathcal{C}\sum_{L=L^{\ast}+1}^{n}\sigma_{L},$
@@
For $\eta=\varepsilon=0$ this is bounded by $O_{L}(L^{-2}2^{-\delta^{2}L}2^{\mathcal{C}_{1}L^{2/3}})\leq O_{L}(2^{-\varepsilon^{\prime}L})$ for some $\varepsilon^{\prime}>0$. Since the expression is continuous in both $\eta$ and $\varepsilon$, we can choose them so small that $\mathbb{E}[\sigma_{L}]\leq O_{L}(2^{-\varepsilon^{\prime}/2L})$. By Markov’s inequality and Borel-Cantelli, this implies that for all but finitely many $L$ we have $\sigma_{L}\leq 2^{-\varepsilon^{\prime}/4L}$. This immediately implies Eq.˜3.32, finishing the proof of absolute continuity of $\widetilde{R}_{n,\gamma,1}(t)$ with respect to $\widetilde{\mu}_{\gamma,\widetilde{a}}$.
@@
$J_{\eta,n}^{\prime}:=\sum_{v\in\mathcal{N}_{n}}\mathbbm{1}_{\{t\in\mathcal{T}_{n,\gamma}^{c}\}}\widetilde{\mu}_{\gamma,\widetilde{a}}[I_{n}(v)]e^{\gamma(1+\eta)(\widetilde{S}_{n,\widetilde{g}}(v)-\widetilde{S}_{n,\widetilde{a}}(v))}\quad\text{$\mathbb{P}$-almost surely.}$ (3.36)
Similar to Eq.˜3.33, we have
@@
Using Eq.˜3.34 again, taking $L:=n$ there, we find some $\mathcal{C}^{\prime}(\eta)\in(0,\infty)$ $\mathbb{P}$-almost surely such that
@@
By taking $\eta,\varepsilon$ close enough to 0 this is summable, which implies Eq.˜3.36. This finishes the proof of uniform integrability of $\widetilde{R}_{n,\gamma}(t)$ with respect to $\widetilde{\mu}_{\gamma,\widetilde{a}}$.